Let L:P1 --> P2 be a function defined by L(p)=tp + 1.

Where p is a linear polynomial in P1.

which of the following is not true?

L is a linear transformation

L is not a linear operator

L is not a linear transformation

L is not a 1-1 function
Answer #1

linear operator. Lip 2 defined by -Cb) = t 5+1 Note that +(2p) = 5.25 +1 = 2t5+1 a L(P) = 2 (tb+1) atpta so LC26) & aL(P). Th